i have a word problem which i cannot find a formula for it? the problem goes like this "some office workers bought a $35 gift for their boss. if there had been two more employees to contribute everyones cost would have been $2 less.How many workers contributed to the gift?
Answer by scott8148(6628)

(35 / n) - 2 = 35 / (n + 2)

multiplying by n(n + 2) ___ 35n + 70 - 2n^2 - 4n = 35n

n^2 + 2n - 35 = 0

factoring ___ (n + 7)(n - 5) = 0

n = 5